OpenStudy (anikate):

the denominator of a fraction is 1 less than twice the numerato. If 7 is added to both the numerator and denominator, the resulting fraction has a value of 7/10. What is the "original" fraction.

OpenStudy (anikate):

help me!

OpenStudy (anikate):

anyone here?

OpenStudy (anonymous):

Similar to the last one. you have x/2x-1 where x is the numerator. Adding 7 to top and bottom gives you x+7/2x-1+7 whic is x+7/2x+6. set it equal to 7/10 and cross multiply to get 10x+70=14x+42. Simplifying gives you 4x=28, so x=7. Plug it back in to get the orignial fraction. 7/2(7)+1=7/15. Check my math but that's how you do it.

OpenStudy (anonymous):

http://www.algebra.com/algebra/homework/Rational-functions/Rational-functions.faq.question.557410.html you can relate to this one @Anikate hope it may help. just change the data.